7. Find the coordinates of a point A, where AB is the diameter of circle whose centre is (2, – 3) and B is (1, 4).

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the coordinates of point A be (x, y). Mid-point of AB is (2, – 3), which is the centre of the circle. Coordinate of B = (1, 4) (2, -3) =((x+1)/2 , (y+4)/2) (x+1)/2 = 2 and (y+4)/2 = -3 x + 1 = 4 and y + 4 = -6 x = 3 and y = -10 The coordinates of A(3,-10).